# Huan Zhang

### v1

@@ 2 The continuous-time dynamical system
$f(x(t))-f(x^{*})=\mathcal{O}\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty.$
@@
In the special case for System (4) when $p=q=1$. This implies that $\mu(t)=1$ and $\tau(t)=t$. System (4) collapses to the following open-loop system:
@@
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
@@
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition.
@@
The system (26) can be as a special case of the system proposed in hetian2025 . Clearly, the parameters of the system (26) satisfy the Assumption 2.1 in hetian2025 . Thus, by Theorem 2.3 in hetian2025 , we have
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
@@
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
@@ 3 The discrete accelerated autonomous primal-dual algorithm
In this section, we propose an accelerated autonomous primal-dual algorithm with adaptive step size defined via gradient and analyze the convergence properties of this algorithm. For simplicity, we restrict our discussion to the case where $q=1$, which gives $\dot{\tau}(t)=\mu(t)$. Obviously, the system (24) can be reformulated as:
@@
For the system (27), we consider the time step size $1$ and set $y(k)\approx y_{k}$, $x(k)\approx x_{k}$, $\tau(k)\approx\tau_{k}$, $\dot{\tau}(k)\approx\gamma_{k}$, $\lambda(k)\approx\lambda_{k}$, and $\mu(k)\approx\mu_{k}$. Evaluating the terms $\frac{d}{dt}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)$ and $\ddot{\tau}(t)$ as the time derivative of $\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)$ and $\dot{\tau}(t)$, respectively. Now, by the time discretization of the system (27), we have
@@
Substituting the first equation and the third equation into the second equation in (28), we have
@@
Based on the above analysis, we are now in the position to introduce the following accelerated autonomous primal-dual algorithm for solving Problem (3).
@@
(attouchbot2025, , Lemma 4.2) Let $\{\mu_{k}\}_{k\geq 0}$ be a positive sequence and $\{\tau_{k}\}_{k\geq 1}$ such that $\tau_{1}=0$ and $\tau_{k}=\sum_{i=0}^{k-2}\mu_{i}$ for all $k\geq 2$. Suppose that there exist $C_{4}>0$ and $a,b,c\geq 0$ such that $b+c>a$ and
@@
Then, there exists $C_{5}>0$ such that for every $k\geq 2$, it holds $\tau_{k+1}\geq C_{5}k^{\frac{b+c+1}{b+c-a}}.$
@@
Now, we prove the convergence rates of the primal-dual gap, the feasibility violation, and the objective residual generated by AAPDA.
@@
Let $\{(x_{k},\lambda_{k})\}_{k\geq 0}$ be the sequence generated by $\rm{AAPDA}$. Suppose that $\{\mu_{k}\}_{k\geq 0}$ is a non-decreasing sequence satisfying $\mu_{k}\geq 1$ for all $k\geq 0$. Also, let $\{\tau_{k}\}_{k\geq 1}$ be a sequence such that $\tau_{1}=0$ and $\tau_{k}=\sum_{i=0}^{k-2}\mu_{i}$ for all $k\geq 2$. Then, for any $\left(x^{*},\lambda^{*}\right)\in\mathcal{S}$, we have
$\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
$\|Ax_{k}-b\|=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
$|f(x_{k})-f(x^{*})|=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
@@
For $\left(x^{*},\lambda^{*}\right)\in\mathcal{S}$, we introduce the energy sequence which defined as
@@
From (35) and the convexity of $\mathcal{L}(\cdot,\lambda^{*})$, we have
@@
$\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{\tau_{k+1}}\right),\textup{ as }k\rightarrow+\infty,$ (42)
@@
For notation simplicity, denote a sequence $\{g_{k}\}_{k\geq 0}$ by
@@
This together with the boundedness of $\{\lambda_{k}\}_{k\geq 0}$ follows that $\sup_{k\geq 0}\|g_{k}\|<+\infty$. Then,
$\|Ax_{k}-b\|\leq\mathcal{O}\left(\frac{1}{\tau_{k+1}}\right),\textup{ as }k\rightarrow+\infty.$ (43)
Combining (42), (43) and the definition of $\mathcal{L}$, we have
$\left|f(x_{k})-f(x^{*})\right|\leq\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})+\|\lambda^{*}\|\|Ax_{k}-b\|\leq\mathcal{O}\left(\frac{1}{\tau_{k+1}}\right),\textup{ as }k\rightarrow+\infty.$ (44)
@@
Then, by using Lemma 2, there exists $C_{3}>0$ such that
@@
This together with (42), (43) and (44) follows that
$\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
$\|Ax_{k}-b\|=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
@@
$|f(x_{k})-f(x^{*})|=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty.$
@@ 4 Numerical experiments
In this section, we validate the theoretical findings of previous sections through numerical experiments. In these experiments, all codes are implemented in MATLAB R2021b and executed on a PC equipped with a 2.40GHz Intel Core i5-1135G7 processor and 16GB of RAM.
@@
Consider the following constrained minimization problem
@@
Figure 1: Numerical results of AAPDA, FPDA and AALM under different dimensions.
@@
In the following experiments, we compare AAPDA with the fast primal-dual algorithm (FPDA) proposed in hehu2022 and the accelerated linearized augmented Lagrangian method (AALM) proposed in xu2017 . Here are parameter settings of algorithms:
@@
As shown in Figure 1, AAPDA achieves significantly superior performance over FPDA and AALM under all three different dimensions. It is worth noting that the convergence speed and accuracy of AAPDA improve with an increase in the value of $p$, and AAPDA can achieve exponential convergence. Moreover, compared to FPDA, Hessian-driven damping term in AAPDA can induces significant attenuation of the oscillations.
@@
Consider the following non-negative least squares problem
@@
where $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$. We generate a random matrix $A\in\mathbb{R}^{m\times n}$ with density $s\in(0,1]$ and a random $b\in\mathbb{R}^{m}$. The nonzero entries of $A$ are independently generated from a uniform distribution in $[0,0.1]$. The original solution $x^{*}$ is obtained via $x^{*}:=A\setminus b$ in MATLAB.
@@
or the number of iterations exceeds 200. For $s=0.5$ and $s=1$, we respectively consider $\theta:=\{10^{-6},10^{-8},10^{-10}\}$.
In the following experiments, we compare AAPDA with FISTA proposed in beck2009 , the proximal inertial algorithm (PIA) proposed in (attouchbot2025, , Algorithm 4) and the accelerated forward-backward method (AFBM) proposed in a2016 . Here are the parameter settings of algorithms:
@@
Figure 2: Numerical results of AAPDA, FISTA, PIA and AFBM when $\theta=10^{-6}$.
@@
Figure 3: Numerical results of AAPDA, FISTA, PIA and AFBM when $\theta=10^{-8}$.
@@
Figure 4: Numerical results of AAPDA, FISTA, PIA and AFBM when $\theta=10^{-10}$.
As shown in Figures 2, 3 and 4, AAPDA demonstrates superior convergence performance compared to FISTA, PIA and AFBM. Specifically, AAPDA achieves a significantly faster convergence rate and consistently reaches a much higher final accuracy across different values of $\theta$. Moreover, AAPDA exhibits more stable performance, while the convergence speed and final accuracy of FISTA, PIA and AFBM are obviously affected by $\theta$.
@@ 5 Conclusion
In this paper, for solving a convex optimization problem with linear equality constraints (3), we propose a primal-dual dynamical system (4), and obtain the fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along the trajectory of System (4). Specifically, the damping in System (4) is a feedback control of the gradient of the Lagrangian function for Problem (3). We also propose an accelerated autonomous primal-dual algorithm (AAPDA), which is derived from the time discretization of System (4), and establish convergence rates for the primal-dual gap, the feasibility violation, and the objective residual.
As shown in this paper, the convergence of the trajectories generated by the primal-dual dynamical system with closed-loop control is still a challenge. To address this, future work will reset the primal-dual dynamical system with closed-loop control and the energy function to achieve the convergence of the trajectories generated by systems.

### v2

@@ 2 The continuous-time dynamical system
$f(x(t))-f(x^{*})=O\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty.$
@@
In the special case of System (4) where $p=q=1$, we have $\mu(t)=1$ and $\tau(t)=t$. Then, System (4) collapses to the following open-loop system:
@@
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=O\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=O\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=O\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
@@
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$, and the upper bound depends only on the initial condition.
@@
System (25) can be seen as a special case of the system proposed in hetian2025 . Clearly, the parameters of System (25) satisfy the Assumption 2.1 in hetian2025 . Thus, by Theorem 2.3 in hetian2025 , we have
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=O\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=O\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
@@
$|f(x(t))-f(x^{*})|=O\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
@@ 3 The discrete accelerated autonomous primal-dual algorithm
In this section, we propose an accelerated autonomous primal-dual algorithm with adaptive step size defined via gradient and analyze the convergence properties of this algorithm. For simplicity, we restrict our discussion to the case where $q=1$, which gives $\dot{\tau}(t)=\mu(t)$. Obviously, System (23) can be reformulated as:
@@
For System (26), we consider a time step size of $1$ and set $y(k)\approx y_{k}$, $x(k)\approx x_{k}$, $\tau(k)\approx\tau_{k}$, $\dot{\tau}(k)\approx\gamma_{k}$, $\lambda(k)\approx\lambda_{k}$, and $\mu(k)\approx\mu_{k}$. We evaluate the terms $\frac{d}{dt}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)$ and $\ddot{\tau}(t)$, which are the time derivative of $\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)$ and $\dot{\tau}(t)$, respectively. Now, by time discretization of System (26), we obtain
@@
Substituting the first equation and the third equation into the second equation in System (27), we have
@@
Based on the above analysis, we are now in a position to introduce the following accelerated autonomous primal-dual algorithm for solving Problem (3).
@@
(attouchbot2025, , Lemma 4.2) Let $\{\mu_{k}\}_{k\geq 0}$ be a positive sequence and let $\{\tau_{k}\}_{k\geq 1}$ be a sequence such that $\tau_{1}=0$ and $\tau_{k}=\sum_{i=0}^{k-2}\mu_{i}$ for all $k\geq 2$. Suppose that there exist constants $C_{4}>0$ and $a,b,c\geq 0$ such that $b+c>a$ and
@@
Then, there exists a constant $C_{5}>0$ such that for any $k\geq 2$, it holds that $\tau_{k+1}\geq C_{5}k^{\frac{b+c+1}{b+c-a}}.$
@@
We now establish the convergence rates of the primal-dual gap, the feasibility violation, and the objective residual generated by AAPDA.
@@
Let $\{(x_{k},\lambda_{k})\}_{k\geq 0}$ be the sequence generated by $\rm{AAPDA}$. Suppose that $\{\mu_{k}\}_{k\geq 0}$ is a non-decreasing sequence satisfying $\mu_{k}\geq 1$ for all $k\geq 0$. Also, let $\{\tau_{k}\}_{k\geq 1}$ be a sequence such that $\tau_{1}=0$ and $\tau_{k}=\sum_{i=0}^{k-2}\mu_{i}$ for all $k\geq 2$. Then, for any $\left(x^{*},\lambda^{*}\right)\in\mathcal{S}$, it holds that
$\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=O\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
$\|Ax_{k}-b\|=O\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
$|f(x_{k})-f(x^{*})|=O\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
@@
For $\left(x^{*},\lambda^{*}\right)\in\mathcal{S}$, we introduce the following energy sequence
@@
From (34) and the convexity of $\mathcal{L}(\cdot,\lambda^{*})$, we obtain
@@
$\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=O\left(\frac{1}{\tau_{k+1}}\right),\textup{ as }k\rightarrow+\infty,$ (41)
@@
For notational simplicity, denote a sequence $\{g_{k}\}_{k\geq 0}$ by
@@
From this, together with the boundedness of $\{\lambda_{k}\}_{k\geq 0}$, it follows that $\sup_{k\geq 0}\|g_{k}\|<+\infty$. Then,
$\|Ax_{k}-b\|\leq O\left(\frac{1}{\tau_{k+1}}\right),\textup{ as }k\rightarrow+\infty.$ (42)
Based on (41), (42) and the definition of $\mathcal{L}$, we obtain
$\left|f(x_{k})-f(x^{*})\right|\leq\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})+\|\lambda^{*}\|\|Ax_{k}-b\|\leq O\left(\frac{1}{\tau_{k+1}}\right),\textup{ as }k\rightarrow+\infty.$ (43)
@@
Thus, by using Lemma 2, it follows that there exists a constant $C_{3}>0$ such that
@@
This together with (41), (42) and (43) gives
$\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=O\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
$\|Ax_{k}-b\|=O\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
@@
$|f(x_{k})-f(x^{*})|=O\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty.$
@@ 4 Numerical experiments
In this section, we validate the theoretical findings of previous sections through numerical experiments. In these experiments, all code is implemented in MATLAB R2021b and executed on a PC equipped with a 2.40GHz Intel Core i5-1135G7 processor and 16GB of RAM.
@@
Consider the following constrained minimization problem:
@@
Figure 1: Numerical results obtained using AAPDA, FPDA and AALM under different dimensions.
@@
In the following experiments, we compare AAPDA with the fast primal-dual algorithm (FPDA) proposed in hehu2022 and the accelerated linearized augmented Lagrangian method (AALM) proposed in xu2017 . Here are the parameter settings for the algorithms:
@@
As shown in Figure 1, AAPDA achieves significantly superior performance over FPDA and AALM under all three different dimensions. It is worth noting that the convergence speed and accuracy of AAPDA improve as the value of $p$ increases, and that AAPDA can achieve exponential convergence. Moreover, compared to FPDA, the Hessian-driven damping term in AAPDA can induces significant attenuation of the oscillations.
@@
Consider the following non-negative least squares problem:
@@
where $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$. We generate a random matrix $A\in\mathbb{R}^{m\times n}$ with density $s\in(0,1]$ and a random vector $b\in\mathbb{R}^{m}$. The nonzero entries of $A$ are independently generated from a uniform distribution on $[0,0.1]$. The true solution $x^{*}$ is obtained via $x^{*}:=A\setminus b$ in MATLAB.
@@
or the number of iterations exceeds 200. For $s=0.5$ and $s=1$, we consider $\theta:=\{10^{-6},10^{-8},\\ 10^{-10}\}$, respectively.
In the following experiments, we compare AAPDA with FISTA proposed in beck2009 , the proximal inertial algorithm (PIA) proposed in (attouchbot2025, , Algorithm 4) and the accelerated forward-backward method (AFBM) proposed in a2016 . Here are the parameter settings for the algorithms:
@@
Figure 2: Numerical results obtained using AAPDA, FISTA, PIA and AFBM when $\theta=10^{-6}$.
@@
Figure 3: Numerical results obtained using AAPDA, FISTA, PIA and AFBM when $\theta=10^{-8}$.
@@
Figure 4: Numerical results obtained using AAPDA, FISTA, PIA and AFBM when $\theta=10^{-10}$.
As shown in Figures 2, 3 and 4, AAPDA demonstrates superior convergence performance compared to FISTA, PIA and AFBM. Specifically, AAPDA achieves a significantly faster convergence rate and consistently reaches a much higher final accuracy across different values of $\theta$. Moreover, AAPDA exhibits more stable performance, while the convergence speed and final accuracy of FISTA and AFBM are affected by $\theta$.
@@ 5 Conclusion
In this paper, for solving a convex optimization problem with linear equality constraints (3), we propose a primal-dual dynamical system (4), and establish fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along its trajectory. A distinctive feature of System (4) is its damping mechanism, which acts as a feedback control driven by the gradient of the Lagrangian function in Problem (3). Building on this continuous-time framework, we introduced an accelerated autonomous primal-dual algorithm (AAPDA) through time discretization of System (4), and derived analogous convergence rates for the primal-dual gap, the feasibility violation, and the objective residual.
Despite these advances, the convergence analysis of trajectories generated by primal-dual dynamical system with closed-loop control remains an open challenge. In future work, we aim to address this limitation by reformulating the system and constructing an appropriate energy function, with the goal of rigorously establishing trajectory convergence.